Hi im trying to figure out how to solve equations with the variable on both sides. Ive been sick with the swine flu so i missed that lesson and now i cant seem to get the hang on how to work out problems like this: q+14=8(q+7)
i would be more than happy if you could help me, thanks

Answer by Alan3354(69443)

q+14=8(q+7)
q+14 = 8q+56
14 = 7q+56
-42 = 7q
q = -6
-------
Don't kiss any pigs. Piggy banks are OK.